Can you please answer the questions from chapter 3 review problems (pg. 90-91)questions 2,4,8,10,12. chapter 4 review problems 2,4,6,8,10,12,16 (pg121) chapter 5 review problems 2,4,6,12,14 (pg.141).Chapter3-Question2-find the x- and y- components of vector R, which has a length of 10.0 cm at 60 degrees. Question4-Vector R has a length of 9.00cm at 240.0 degrees. Find its x-and y components. Question8-A hiker is plotting his course on a map with a scale of 1.00cm=3.00km. If the hiker walks 2.50 cm north, then turns south and walks 1.50 cm, what is the actual displacement of the hiker in Km? Question10- A co-pilot is charting her course on a map with a scale of 1.00cm=20.0 Km. If the plane is charted to head 25.0 degrees north of east for 16.0cm, north for 6.00cm, and west for 5.00cm, what is the actual displacement of the plane in Km? Question12-Vector R has x-component = -5.00 and y-component = +10.0. Find its length. Chapter4, Question 2- A commercial jet flies at 550 mi/h for 3000 mi. For how much time does the jet fly? Question 4, A glider flies south east (at 320.0 degrees) at 25.0 Km/h. a wind blows at 12.0 Km/h from 15.0 degrees south of west. What is the new velocity of the glider with respect to the ground in standard position? Question 6- A race car goes from rest to 150 Km/h with an acceleration of 6.0 m/s squared. How many seconds does it take?. Question8- A skateboard starts from rest and accelerates at a rate of 1.30 m/s squared for 3.00s What is his final velocity? Question 10 A train has a final final velocity of 110 Km/h . It accelerated for 36 s at 0.50 m/s squared. What was its initial velocity? Question 12-A truck accelerates from rest to 120Km/h in 13 s. Find its acceleration. Question 16- A rock is thrown down with an initial speed of 10.0 m/s from a bridge to the water below. It takes 2.75 s for the rock to hit the water. (a) Find the speed of the rock as it hits the water. (b) How high is the bridge above the water? Chapter5. Question2- An 825-N force is required to pedal a bike with an acceleration of 11.0 m/s squared . What is the mass of the bike and rider? Question4- What is the force necessary for a 2400-Kg truck to accelerate at a rate of 8.0 m/s squared? Question 6- A 340-N box has a frictional force of 57  N. Find the coefficient of Kinetic friction. Question12- Find the total force necessary to give a 280-Kg motorcycle an acceleration of 3.20m/s squared. Question14- A crated garden tractor weighs 375 N. What force is needed to start the crate sliding on a wooden floor when the coefficient of static friction is 0.40?         
Answer #1

Ans:-

Chapter3, Question2

Given data:- R= 10cm, θ=60deg

X= Rcos θ = 10*cos60 =5cm

Y = Rsin θ =10sin60 =8.66cm

Question 4

R= 9cm θ= 240deg

X = Rcos θ = 9*cos240 = -4.5cm

Y = Rsin θ = 9*sin240 = -7.79cm

Question8

1cm = 3km,2.50cm N, 1.50cm S

D = Xf –Xi

Xi = 0 here initial position is zero

D = (2.50-1.50) – 0 = 1cm

Means 3Km

Question 10

1cm = 20Km,A= 16cm θ= 25deg, y= 6cm X= -5cm

Ax= 16cos25 = 14.5cm

Ay= 16sin25 = 6.76cm

X1 = Ax + X = 14.5 – 5 =9.5cm

Y1 = Ay+ y = 6.76+6 = 12.76cm

D = sqrt(9.5^2 +12.76^2) =15.9cm

Means 318.16Km

Question12

Rx= -5,Ry= 10

R= sqrt(Rx^2 +Ry^2)

= sqrt(-5^2 +10^2 )

= 11.18unit